OpenStudy (anonymous):

How do I express 4^-2 without exponents?

OpenStudy (agent0smith):

For exponents, \[\large a ^{-b} = \frac{ 1 }{ a^b }\] so \[\large 4^{-2}=\frac{ 1 }{ 4^2 }\] can you simplify it from here?

OpenStudy (anonymous):

music is my life the lyrics are my story -newr3d

OpenStudy (anonymous):

1/16?

OpenStudy (anonymous):

correct! :D